A gas at a temperature of 127 0C is heated in a freely expandable container until its volume doubles. What is the new temperature of the gas?  


1
Expert's answer
2022-04-06T03:07:03-0400

T1 = 127 C = 400 K;

V2 = 2 * V1;

V1/T1 = V2/T2 or V1/T1 = 2 * V1/T2;

T2 * V1 = 2 * V1 * T1;

T2 = 2 * V1 * T1/V1 = 2 * T1 = 2 * 400 = 800 K = 527 C.

Answer: 527 C.